The product of 3/5 and 3/4 is


1
Expert's answer
2021-05-07T10:11:35-0400

Multiply the numerators and denominators

numerator product "=3\times 3=9"

denominator product "=5\times 4=20"

Therefore;


"\frac{3}{5}\times \frac{3}{4}=\frac{9}{20}"


In decimal form

"=0.45"